OpenStudy (anonymous):

please can someone help me with solution to measure theory and integration question???

OpenStudy (anonymous):

find the length of the set \[\cup_{k=1}^{\infty} \left\{ x:\frac{ 1 }{ k + 1 } \le x \le \frac{ 1 }{ k } \right\}\]

OpenStudy (anonymous):

please help

OpenStudy (tboi):

What is the first step do you think you should do?

OpenStudy (anonymous):

i don"t know

OpenStudy (tboi):

Have you done questions like this before?

OpenStudy (tboi):

@GIL.ojei

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@tboi

OpenStudy (tboi):

So, what I would is look on the internet like google for examples to questions like that, ok?

OpenStudy (tboi):

@GIL.ojei

OpenStudy (anonymous):

please solve this first, that will also give me a clue

OpenStudy (tboi):

Sorry it is against the code of conduct to give out direct answers.

OpenStudy (anonymous):

you are not giving out answers, you are giving a prove . which includes statements and directions too

OpenStudy (tboi):

I know but I cant do that I have to let you do it and find the answer your self, I am so sorry but all i can do is guide you toward the answer.

OpenStudy (anonymous):

ok. please guide me . what should i do?

OpenStudy (tboi):

Ok first what I would do is simplify the equation to make it easier to solve. Tell me when you are done doing that.

OpenStudy (anonymous):

simplifing, i got k ≤ -1 and k ≤ 1/x a few moments ago

OpenStudy (tboi):

@mathmale here he can help ok I have to go I hope you get what you need.

OpenStudy (mathmale):

@GIL.ojei: My first suggestion would be that y ou let k=1 and write the resulting inequality, then k=2, do the same, for k=1 to 5 or 6. These concrete examples should give you an idea of what the set in question looks like. Please share any work you do.

OpenStudy (anonymous):

ok. i will post my result soon. thank you very much

OpenStudy (anonymous):

For k=1 in the function I have , 1/2 <=x<=1 and for k=2, I have 1/3<=x<=1/2 and it continues that manner. So what will be my conclusion sir?

OpenStudy (anonymous):

What next should I do or how should I conclude the the length?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

@ganeshie8